In printing an article of 48,000 words, a printer decides to use two sizes of type. Using the larger type, a printed page contains 1,800 words. Using smaller type, a page contains 2,400 words. The article is allotted 21 full pages in a magazine. How many pages must be in smaller type?
Answer #1
Let x = pages with 1800 words and y = pages with 2400 words.

From your information:

1800x + 2400y = 48,000

x + y = 21

Then y = 21 - x

Substitute (21-x) for y in the first equation and solve for y. Put the value of y in the second equation to get x. To check, put both values into the first equation.

I'll let you do the calculations.
answered by: leeann
Answer #2

small print pages: x

large print pages: 21-x


2400x + 1800(21-x) = 48000


solve for x

===

600x = 48000-37800

x = 17

answered by: mathboy